If two adjacent edges of a given square were made to be 2 cm longer, the area of the resulting square contains 84 more square centimeters than the original. How long is an edge of the original square?
Answer by mananth(16946)


Let the length of square be x cm
the increased length = x+2
x^2+84 = (x+2)^2
x^2+84=x^2+4x+4
4x=80
x=20 cm
the length of original square